John is 5 years older now than twice Jills age last year. Their current age now is 59. What is both their ages now?
Answer by fractalier(6550)

Call John and Jill's ages NOW, x and y. Then we write
x = 2(y - 1) + 5 and (I think you mean)
x + y = 59
Now substitute and get
2(y-1)+5 + y = 59
2y - 2 + 5 + y = 59
3y + 3 = 59
3y = 56
y = 56/3 = 18 2/3
x = 40 1/3